OpenStudy (anonymous):

PLEASES HELP ME!!! Give my like 5 minutes to write the question in the answer box!

OpenStudy (anonymous):



OpenStudy (anonymous):

to simplify it...factorise dem...so dat simillar terms wil get cancel

OpenStudy (anonymous):

I tried but Im not getting the correct answer

OpenStudy (anonymous):

5x2-8x+3 =5x2- 5x - 3x + 3 =5x (x - 1) -3 (x - 1) =(5x - 3) (x - 1) x2+4x-5 =(x+5)(x-1) 20x-12 =4(5x-3) x2-6x-55 =(x-11)(x+5) put all these values in above eqn simillar terms wil get cancel ... 4/(x-11)